Let z be a standard normal random variable.

P(0 < z < 0.41)=______ .

For the value a =______, P(z > a) = 0.2296
Answer #1

Solution,

Using standard normal table,  

P( 0 < Z < 0.41)

= P( Z < 0.41) - P( Z < 0)

= 0.6591 - 0.5

= 0.1591

Using standard normal table,

P(Z > a) = 0.2296

= 1 - P(Z < a) = 0.2296

= P(Z < a) = 1 - 0.2296

= P(Z < a ) = 0.7704

= P(Z < 0.74 ) = 0.  

a = 0.74